# Properties preserved by classes of Chu transforms

### v1

@@ 4. Preservation of compactness and inconsistency-flow formulas
$\varphi=\left\{\alpha(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})\mathchoice{\>}{\>}{\,}{\,}\middle|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha\text{ is }\mathcal{L}^{\mathrm{II}}_{\omega,\omega}\text{-inconsistency-flow and }T\vdash\forall\bm{v}^{\mathsf{p}}\forall\bm{v}^{\mathsf{s}}(\varphi\rightarrow\alpha)\right\}.$
We intend to prove that taking $\theta=\bigwedge F$ for some finite $F\subseteq\varphi$ will satisfy the theorem. We first show that
(eq-6) $T\vdash\forall\bm{v}^{\mathsf{p}}\forall\bm{v}^{\mathsf{s}}\Bigl{(}\bigwedge\varphi\rightarrow\psi\Bigr{)}.$
If $T$ is inconsistent then we are done, otherwise let $\langle Y,s,B\rangle$ be a model of $T$ and let $\bm{y}\in Y$ and $\bm{b}\in B$ be such that $\langle Y,s,B\rangle\models\alpha[\bm{y},\bm{b}]$ for each $\alpha\in\varphi$. We need to prove that $\langle Y,s,B\rangle\models\psi[\bm{y},\bm{b}]$.
@@
$\Sigma=\left\{\varphi\land\lnot\gamma\mathchoice{\>}{\>}{\,}{\,}\middle|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\gamma(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})\text{ is }\mathcal{L}^{\mathrm{II}}_{\omega,\omega}\text{-inconsistency-flow and }\langle Y,s,B\rangle\models\lnot\gamma[\bm{y},\bm{b}]\right\}.$
@@
[…]
is consistent, for otherwise we would have $T\vdash\forall\bm{v}^{\mathsf{p}}\forall\bm{v}^{\mathsf{s}}\bigl{(}\varphi\rightarrow(\gamma_{1}\lor\dots\lor\gamma_{m})\bigr{)}$ and therefore $(\gamma_{1}\lor\dots\lor\gamma_{m})\in\varphi$, a contradiction. Since each of these sets is consistent, by compactness of two-sorted first-order logic the whole set is consistent.
Hence, there exists a model $\langle X,r,A\rangle$ of $T$ and $\bm{x}\in X$ and $\bm{a}\in A$ such that each of the following holds:
@@
Now Lemma 4.11 gives two elementary extensions
$\langle X,r,A\rangle\preceq\langle X^{\prime},r^{\prime},A^{\prime}\rangle\mbox{ and }\langle Y,s,B\rangle\preceq\langle Y^{\prime},s^{\prime},B^{\prime}\rangle$
as well as a dense Chu transform $(f,g)$ from $\langle X^{\prime},r^{\prime},A^{\prime}\rangle$ to $\langle Y^{\prime},s^{\prime},B^{\prime}\rangle$ such that $f(\bm{x})=\bm{y}$ and $g(\bm{b})=\bm{a}$. From this, we conclude that $\langle Y,s,B\rangle\models\psi[\bm{y},\bm{b}]$, as we wanted to show.
This completes the proof of (eq-6). By compactness again, there exists a finite subset $F\subseteq\varphi$ such that $T\vdash\forall\bm{v}^{\mathsf{p}}\forall\bm{v}^{\mathsf{s}}\Bigl{(}\bigwedge F\rightarrow\psi\Bigr{)}$. Then it is sufficient to take $\theta$ to be $\bigwedge F$.
@@ 5. Ultrafilters as Chu spaces
We shall be coding ultrafilters as partially ordered sets, so we start with a known characterization of Chu transforms in this context. First of all, a partially ordered set $\bigl{\langle}P,\leq^{P}\bigr{\rangle}$ can be naturally represented as a Chu space $\bigl{\langle}P,\leq^{P},P\bigr{\rangle}$. Chu transforms between such Chu spaces are characterized syntactically by the following proposition due to Schmidt.
Proposition 5.1 (Schmidt [26, Section 8]).
@@
5.1. Chu transforms correspond to the Rudin-Keisler order
An ultrafilter $U$ over a set $I$ is considered as a partially ordered set $\langle U,\supseteq\rangle$. We prove that, in this framework, Chu transforms correspond to an ordering of ultrafilters introduced in the sixties independently by Mary Ellen Rudin and Keisler. Let us recall the definition:
@@ 5.1. Chu transforms correspond to the Rudin-Keisler order
Let us define a function $h\colon Y\to\kappa$ as follows: for all $y\in Y$
@@
We claim that $h$ witnesses the Rudin-Keisler reducibility of $U$ to $V$.
@@
For the first part of the claim, let $y\in f(X)\cap Y$, we want to show that $h(y)\in X$. Suppose not; then $X\subseteq\kappa\setminus\{h(y)\}$. By condition (1) of Proposition 5.1, the function $f$ is inclusion-preserving, which gives $f(X)\subseteq f(\kappa\setminus\{h(y)\})$. But $y\in f(X)$ and therefore $y\in f(\kappa\setminus\{h(y)\})$, contradicting the definition of $h$. ∎
@@
$\begin{split}f\colon U&\longrightarrow V\\ X&\longmapsto h^{-1}[X]\end{split}\qquad\quad\text{and}\qquad\quad\begin{split}g\colon V&\longrightarrow U\\ Y&\longmapsto h[Y]\end{split}$
@@
5.2. Closure under reduced products
Given an abstract logic $\langle M,\models,S\rangle$ and a filter $F$ over a set $I$, the $F$-closure of $\langle M,\models,S\rangle$ is the abstract logic $\bigl{\langle}\prescript{I}{}{M},\models_{F},S\bigr{\rangle}$, where the relation $\models_{F}$ is defined as
$\langle\mathfrak{M}_{i}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}i\in I\rangle\models_{F}\sigma\iff\{i\in I\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\mathfrak{M}_{i}\models\sigma\}\in F.$
The $F$-closure can be thought as the closure under reduced products modulo $F$. An alternative approach, which consists of looking at the fragment of $S$ which is invariant under a given ultraproduct operation, is undertaken by Joseph Sgro [27].
We begin by observing that the $F$-closure always yields a sublogic of the original one.
Lemma 5.5.
Let $\langle M,\models,S\rangle$ be an abstract logic and let $F$ be a filter over a set $I$. Then $\bigl{\langle}\prescript{I}{}{M},\models_{F},S\bigr{\rangle}\leq\langle M,\models,S\rangle$.
@@ 5.2. Closure under reduced products
To find a dense Chu transform from $\langle M,\models,S\rangle$ to $\bigl{\langle}\prescript{I}{}{M},\models_{F},S\bigr{\rangle}$, let $f\colon M\to\prescript{I}{}{M}$ be defined as follows: if $\mathfrak{M}\in M$, then $f(\mathfrak{M})$ is the constant sequence $\langle\mathfrak{M}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}i\in I\rangle$. Let $g\colon S\to S$ be the identity function. Then clearly for each $\mathfrak{M}\in M$ and $\sigma\in S$ we have $\mathfrak{M}\models\sigma\iff f(\mathfrak{M})\models_{F}\sigma$, therefore the pair $(f,g)$ is a Chu transform.
To show that $(f,g)$ is dense, let $\sigma\in S$ be such that there exists $\langle\mathfrak{M}_{i}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}i\in I\rangle$ such that $\langle\mathfrak{M}_{i}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}i\in I\rangle\models_{F}\sigma$. By definition, there exists in particular some $i\in I$ such that $\mathfrak{M}_{i}\models\sigma$, hence $f(\mathfrak{M}_{i})\models_{F}\sigma$. ∎
Furthermore, this operation preserves both the sublogic relation and the Rudin-Keisler ordering.
Proposition 5.6.
Let $\langle M,\models,S\rangle$ and $\langle M^{\prime},\models^{\prime},S^{\prime}\rangle$ be two abstract logics, let $U$ be an ultrafilter over a set $I$ and $V$ be an ultrafilter over a set $J$. If $\langle M^{\prime},\models^{\prime},S^{\prime}\rangle\leq\langle M,\models,S\rangle$ and $U\leq_{\mathrm{RK}}V$ then
$\bigl{\langle}\prescript{J}{}{M^{\prime}},\models^{\prime}_{V},S^{\prime}\bigr{\rangle}\leq\bigl{\langle}\prescript{I}{}{M},\models_{U},S\bigr{\rangle}$
Proof.
Let $(f,g)$ be a dense Chu transform from $\langle M,\models,S\rangle$ to $\langle M^{\prime},\models^{\prime},S^{\prime}\rangle$. Let $h\colon J\to I$ be a function such that for each $X\subseteq I$, $X\in U\iff h^{-1}[X]\in V$.
We define $\bar{f}\colon\prescript{I}{}{M}\to{{}^{J}}M^{\prime}$ as follows: given $\langle\mathfrak{M}_{i}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}i\in I\rangle\in\prescript{I}{}{M}$, let
$\bar{f}(\langle\mathfrak{M}_{i}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}i\in I\rangle)=\left\langle f(\mathfrak{M}_{h(j)})\mathchoice{\>}{\>}{\,}{\,}\middle|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}j\in J\right\rangle.$
For every $\langle\mathfrak{M}_{i}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}i\in I\rangle\in\prescript{I}{}{M}$ and $\sigma^{\prime}\in S^{\prime}$, we have
$h^{-1}[\{i\in I\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\mathfrak{M}_{i}\models g(\sigma^{\prime})\}]=\left\{j\in J\mathchoice{\>}{\>}{\,}{\,}\middle|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\mathfrak{M}_{h(j)}\models g(\sigma^{\prime})\right\}=\left\{j\in J\mathchoice{\>}{\>}{\,}{\,}\middle|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}f(\mathfrak{M}_{h(j)})\models^{\prime}\sigma^{\prime}\right\}$
and therefore
$\{i\in I\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\mathfrak{M}_{i}\models g(\sigma^{\prime})\}\in U\iff\left\{j\in J\mathchoice{\>}{\>}{\,}{\,}\middle|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}f(\mathfrak{M}_{h(j)})\models^{\prime}\sigma^{\prime}\right\}\in V,$
which means that $(\bar{f},g)$ is indeed a Chu transform from the Chu space $\bigl{\langle}\prescript{I}{}{M},\models_{U},S\bigr{\rangle}$ to $\bigl{\langle}\prescript{J}{}{M^{\prime}},\models^{\prime}_{V},S^{\prime}\bigr{\rangle}$.
To check the density of $(\bar{f},g)$, let $\sigma^{\prime}\in S^{\prime}$ be such that there exists $\langle\mathfrak{M}^{\prime}_{j}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}j\in J\rangle$ such that $\langle\mathfrak{M}^{\prime}_{j}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}j\in J\rangle\models^{\prime}_{V}\sigma^{\prime}$. Then, in particular, there exists some $j\in J$ such that $\mathfrak{M}^{\prime}_{j}\models^{\prime}\sigma^{\prime}$. By the density of $(f,g)$, there exists $\mathfrak{M}\in M$ such that $f(\mathfrak{M})\models^{\prime}\sigma^{\prime}$. Then the constant sequence $\langle\mathfrak{M}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}i\in I\rangle$ is such that $\bar{f}(\langle\mathfrak{M}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}i\in I\rangle)\models^{\prime}_{V}\sigma^{\prime}$. ∎
Lastly, we discuss when the $F$-closure is $\langle\kappa,\lambda\rangle$-absolutely closed or $\langle\kappa,\lambda\rangle$-compact.
Definition 5.7.
Let $\kappa\leq\lambda$ be infinite cardinals. A filter $F$ over ${[\lambda]}^{<\kappa}$ is fine if for every $\alpha<\lambda$ we have $\left\{Z\in{[\lambda]}^{<\kappa}\mathchoice{\>}{\>}{\,}{\,}\middle|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha\in Z\right\}\in F$.
The next proposition is the abstract logic version of the classic fact (see Keisler and Tarski [16]) that the $\langle\kappa,\lambda\rangle$-compactness of the infinitary logic $\mathcal{L}_{\kappa,\kappa}$ follows from the existence of a ${<}\kappa$-complete fine ultrafilter over ${[\lambda]}^{<\kappa}$.
Proposition 5.8.
Let $\kappa\leq\lambda$ be infinite cardinals:
  1. (1)
    if $F$ is a fine filter over ${[\lambda]}^{<\kappa}$, then the $F$-closure of any abstract logic is $\langle\kappa,\lambda\rangle$-absolutely closed;
  2. (2)
    if $U$ is a ${<}\kappa$-complete fine ultrafilter over ${[\lambda]}^{<\kappa}$, then the $U$-closure of any abstract logic is $\langle\kappa,\lambda\rangle$-compact.
Proof.
(1) Let $F$ be a fine filter over $I={[\lambda]}^{<\kappa}$ and let $\langle M,\models,S\rangle$ be an abstract logic: we have to show that $\bigl{\langle}\prescript{I}{}{M},\models_{F},S\bigr{\rangle}$ is $\langle\kappa,\lambda\rangle$-absolutely closed. To verify (eq-2), let $\{\sigma_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha<\lambda\}\subseteq S$ and suppose that for every $Z\in I$ there exists a consistent $\sigma_{Z}\in S$ such that for every $\mathfrak{M}\in\prescript{I}{}{M}$, if $\mathfrak{M}\models_{F}\sigma_{Z}$ then for every $\alpha\in Z$, $\mathfrak{M}\not\models_{F}\sigma_{\alpha}$. For each $Z\in I$, choose $\mathfrak{M}_{Z}\in M$ such that $\mathfrak{M}_{Z}\models\sigma_{Z}$ and note that for all $\alpha<\lambda$
$\{Z\in I\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\mathfrak{M}_{Z}\not\models\sigma_{\alpha}\}\supseteq\{Z\in I\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha\in Z\}\in F,$
hence $\langle\mathfrak{M}_{Z}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}Z\in I\rangle\not\models_{F}\sigma_{\alpha}$, as desired.
(2) Suppose $U$ is a ${<}\kappa$-complete fine ultrafilter over $I={[\lambda]}^{<\kappa}$ and let $\langle M,\models,S\rangle$ be an abstract logic: we have to show that $\bigl{\langle}\prescript{I}{}{M},\models_{U},S\bigr{\rangle}$ is $\langle\kappa,\lambda\rangle$-compact. To verify (eq-1), let $\{\sigma_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha<\lambda\}\subseteq S$ and suppose that for every $Z\in I$ there exists $\mathfrak{M}\in\prescript{I}{}{M}$ such that for every $\alpha\in Z$, $\mathfrak{M}\not\models_{U}\sigma_{\alpha}$. Since $U$ is a ${<}\kappa$-complete ultrafilter, for each $Z\in I$ it is possible to choose $\mathfrak{M}_{Z}\in M$ such that for every $\alpha\in Z$, $\mathfrak{M}_{Z}\not\models\sigma_{\alpha}$. Then the conclusion follows as in the first point. ∎
A deeper connection between combinatorial properties of ultrafilters and compactness properties of abstract logics is given by the abstract compactness theorem of Johann Makowsky and Saharon Shelah [20, Theorem 3.9].

### v3

@@ 4. Preservation of compactness and inconsistency-flow formulas
$\Phi=\left\{\alpha(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})\mathchoice{\>}{\>}{\,}{\,}\middle|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha\text{ is }\mathcal{L}^{\mathrm{II}}_{\omega,\omega}\text{-inconsistency flow and }T\vdash\forall\bm{v}^{\mathsf{p}}\forall\bm{v}^{\mathsf{s}}(\varphi\rightarrow\alpha)\right\}.$
We intend to prove that taking $\theta=\bigwedge F$ for some finite $F\subseteq\Phi$ will satisfy the theorem. We first show that
(5) $T\vdash\forall\bm{v}^{\mathsf{p}}\forall\bm{v}^{\mathsf{s}}\Bigl(\bigwedge\Phi\rightarrow\psi\Bigr).$
If $T$ is inconsistent, then (5) holds trivially. Otherwise, let $\langle Y,s,B\rangle$ be a model of $T$ and let $\bm{y}\in Y$ and $\bm{b}\in B$ be such that $\langle Y,s,B\rangle\models\alpha[\bm{y},\bm{b}]$ for each $\alpha\in\Phi$; we need to prove that $\langle Y,s,B\rangle\models\psi[\bm{y},\bm{b}]$. By the downward Löwenheim-Skolem theorem, we can assume that $\langle Y,s,B\rangle$ is countable.
@@
$\Sigma=\left\{\varphi\land\lnot\gamma\mathchoice{\>}{\>}{\,}{\,}\middle|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\gamma(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})\text{ is }\mathcal{L}^{\mathrm{II}}_{\omega,\omega}\text{-inconsistency flow and }\langle Y,s,B\rangle\models\lnot\gamma[\bm{y},\bm{b}]\right\}.$
@@
$T\cup\{\varphi(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})\land\lnot\gamma_{i}(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}1\leq i\leq m\}$
is consistent, for otherwise we would have $T\vdash\forall\bm{v}^{\mathsf{p}}\forall\bm{v}^{\mathsf{s}}\bigl(\varphi\rightarrow(\gamma_{1}\lor\dots\lor\gamma_{m})\bigr)$ and therefore $(\gamma_{1}\lor\dots\lor\gamma_{m})\in\Phi$, a contradiction. Since each of these sets is consistent, by compactness of two-sorted first-order logic the whole set is consistent.
Hence, there exists a countable model $\langle X,r,A\rangle$ of $T$ and $\bm{x}\in X$ and $\bm{a}\in A$ such that each of the following holds:
@@
Now Theorem 4.12 applied to the model pair $(\langle X,r,A\rangle,\langle Y,s,B\rangle)$ gives two countable elementary extensions
$X,r,A$ and $Y,s,B\langle$ $\rangle\preceq\langle X^{\prime},r^{\prime},A^{\prime}\rangle\text${ and $\langle$ $\rangle\preceq\langle Y^{\prime},s^{\prime},B^{\prime}\rangle$
such that $(\langle X^{\prime},r^{\prime},A^{\prime}\rangle,\langle Y^{\prime},s^{\prime},B^{\prime}\rangle)$ is recursively saturated as a model pair. By Lemma 4.13 applied to $\langle X^{\prime},r^{\prime},A^{\prime}\rangle$ and $\langle Y^{\prime},s^{\prime},B^{\prime}\rangle$, there exists a dense Chu transform $(f,g)$ from $\langle X^{\prime},r^{\prime},A^{\prime}\rangle$ to $\langle Y^{\prime},s^{\prime},B^{\prime}\rangle$ such that $f(\bm{x})=\bm{y}$ and $g(\bm{b})=\bm{a}$. From this, we conclude that $\langle Y,s,B\rangle\models\psi[\bm{y},\bm{b}]$, as we wanted to show.
This completes the proof of (5). By compactness again, there exists a finite subset $F\subseteq\Phi$ such that $T\vdash\forall\bm{v}^{\mathsf{p}}\forall\bm{v}^{\mathsf{s}}\Bigl(\bigwedge F\rightarrow\psi\Bigr)$. Then it is sufficient to take $\theta$ to be $\bigwedge F$.
@@ 5. Ultrafilters as Chu spaces
We shall be coding ultrafilters as partially ordered sets, so we start with a known characterization of Chu transforms in this context. First of all, a partially ordered set $\bigl\langle P,\leq^{P}\bigr\rangle$ can be represented as a separated and extensional Chu space $\bigl\langle P,{\leq^{P}},P\bigr\rangle$. Chu transforms between such Chu spaces were first considered by Benado [9] and then systematically studied by Jürgen Schmidt [34], who found the following characterization.
Proposition 5.1 (Schmidt [34, §8]).
@@
An ultrafilter $U$ over a set $I$ is considered as a partially ordered set $\langle U,\supseteq\rangle$. We prove that, in this framework, Chu transforms correspond to an ordering of ultrafilters introduced in the sixties independently by Mary Ellen Rudin and Keisler. Let us recall its definition:
@@
Let us define $h\colon Y\to\kappa$ as follows: for all $y\in Y$
@@
We claim that the function $h$ witnesses the Rudin-Keisler reducibility of $U$ to $V$.
@@
For the first part, let $y\in f(X)\cap Y$, we want to show that $h(y)\in X$. Suppose not; then $X\subseteq\kappa\setminus\{h(y)\}$. By condition (1) of Proposition 5.1, the function $f$ is inclusion-preserving, which gives $f(X)\subseteq f(\kappa\setminus\{h(y)\})$. But $y\in f(X)$ and therefore $y\in f(\kappa\setminus\{h(y)\})$, contradicting the definition of $h$. ∎
@@
$\begin{split}f\colon U&\longrightarrow V\\ X&\longmapsto h^{-1}[X]\end{split}\qquad\text{and}\qquad\begin{split}g\colon V&\longrightarrow U\\ Y&\longmapsto h[Y]\end{split}$
@@
Theorem 5.4 generalizes the following well-known result, credited by Rudin to Hewitt Kenyon.
Corollary 5.5 (Rudin [32, Theorem 1.5]).
Let $U$ and $V$ be non-principal ultrafilters over $\omega$. If the partially ordered sets $\langle U,\supseteq\rangle$ and $\langle V,\supseteq\rangle$ are isomorphic, then there exists a bijective function $\pi\colon\omega\to\omega$ such that $V=\{\pi[X]\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}X\in U\}$.
@@
Let $f\colon U\to V$ be an isomorphism of partially ordered sets. Proposition 5.1 implies that $\bigl(f,f^{-1}\bigr)$ is a Chu transform from $\langle U,\supseteq,U\rangle$ to $\langle V,\supseteq,V\rangle$ and $\bigl(f^{-1},f\bigr)$ is a Chu transform from $\langle V,\supseteq,V\rangle$ to $\langle U,\supseteq,U\rangle$. By a double application of Theorem 5.4 we obtain that $U\leq_{\mathrm{RK}}V$ and $V\leq_{\mathrm{RK}}U$, which is equivalent (see, e.g., Rudin [31, III.A]) to the existence of a bijective function $\pi\colon\omega\to\omega$ such that $V=\{\pi[X]\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}X\in U\}$. ∎